HUBBLE SPACE TELESCOPE DAILY REPORT #1421 PERIOD COVERED: 0000Z (UTC) 06/27/95 - 0000Z (UTC) 06/28/95 Daily Status Report as of 179/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6143 (WF/PC-2 Cycle 4 Photometric Calibration: Smaller Set, Part II) The WF/PC-2 was used to make single CCD calibration observations of standard star GRW+70D582 to monitor the photometric stability and quantum efficiency of the PC1 and WF3. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed WF/PC-2 5568 (WF/PC-2 Cycle 4 Calibration: Decontamination) The WF/PC-2 was used to make internal flat field calibration observation using the VIS lamp and the internal source as part of decontamination operations. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 5663 (WF/PC-2 Cycle 4: Photometric Calibration Monitor: Fields Part 2) The WF/PC-2 was used to make observations of globular cluster NGC-5139 with various filters in order to check the photometric calibration of the WF/PC-2. The observations were completed as planned, and no anomalies were reported. 1.4 Completed WF/PC-2 5560 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 1) The WF/PC-2 was used to make several internal bias, K-Spot, and flat calibration observations to monitor the CCD camera electronics, the optical alignment, and to check for contamination. The observations were completed as planned, and no anomalies were noted. 1.5 Completed FOS/RD 6016 (Entrainment and Hotspots in the HH47 Protostar Jet) The Faint Object Spectrograph (Red Side) was used to make observations of Herbig-Haro objects HH47 and HH46. The main purpose is to obtain spectra of the jet in protostar HH47 to investigate the mass exchange between the stream and the surrounding medium, and to look for differences between Mach disk emissions and bow shock emissions. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were reported. 1.6 Completed WF/PC-2 6251 (Medium Deep Survey - Crafting Test) The WF/PC-2 was used in this continuation of the medium deep survey. The proposal is targeting rare objects (lenses, peculiar galaxies, faint stars) and sources related to the faint blue galaxy problem. This observation was completed with no reported problems. 1.7 Completed GHRS/2 5885 (GHRS- Observations of Candidate Beta-Pictoris Circumstellar Gaseous Disks) The High Resolution Spectrograph (Side Two) was used to make observation of stars HD215789 and HD256 which are candidates to have Beta Pictoris-like gaseous disks. The objective is to confirm the existence of circumstellar gas, and to understand its origin. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 4 One acquisition at 179/0722Z failed on FGS #2, and successfully went to coarse track backup on FGS #3. The FHST full maneuver update at 0704Z (in LOS) and an FHST map (2,3) at 0810 had both shown attitude errors. HSTAR 5224 was written. Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 9 Successful : 9 2.3 Significant HSTARs: None 2.4 Operations Notes: No operational problems were reported. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-